# EFX Exists for Three Types of Agents

### v1

@@ 2 Preliminaries
Consider an allocation $X$. Suppose agent $a_{i}$ $g$-champions agent $a_{j}$ and does not $g$-champion $a_{\ell}$. Suppose agent $a_{k}$, for $k\neq i$, $g$-champions agent $a_{\ell}$. The bundle $X_{a_{i}}$ is decomposed as $T_{a_{j}}^{a_{i}}$ and $B_{a_{j}}^{a_{i}}$. Similarly, $X_{a_{\ell}}$ is decomposed into $T_{a_{\ell}}^{a_{k}}$ and $B_{a_{\ell}}^{a_{k}}$. Then $T_{a_{j}}^{a_{i}}>_{a_{i}}T_{a_{\ell}}^{a_{k}}$.
@@
Agent $a_{i}$ $g$-champions $a_{j}$ but not $a_{\ell}$. Therefore, $T_{a_{j}}^{a_{i}}\cup g>_{a_{i}}X_{a_{i}}>_{a_{i}}T_{a_{\ell}}^{a_{k}}\cup g$, which implies $T_{a_{j}}^{a_{i}}>_{a_{i}}T_{a_{\ell}}^{a_{k}}$. ∎
2.1 Three Types of Agents
Given an instance $\langle\mathcal{N},\mathcal{M},\mathcal{V}\rangle$ of discrete fair division, we say that the agents are of three types if the set of agents can be partitioned as $\mathcal{N}=\mathbb{A}\uplus\mathbb{B}\uplus\mathbb{C}$, where agents in each of the sets $\mathbb{A}$, $\mathbb{B}$, and $\mathbb{C}$ have identical valuation functions. We refer to these as the three groups of agents. In particular, we denote the agents in group $\mathbb{A}$ as $\mathbb{A}=\{a_{1},a_{2},a_{3},\dots,a_{p}\}$, all of whom have the same additive valuation function $v_{a}$. Similarly, the agents in group $\mathbb{B}=\{b_{1},b_{2},\dots,b_{q}\}$ and group =$\mathbb{C}=\{c_{1},c_{2},\dots,c_{r}\}$ have identical additive valuation functions $v_{b}$ and $v_{c}$, respectively. Let $p,q$ and $r$ be the number of agents in groups $\mathbb{A},\mathbb{B}$ and $\mathbb{C}$ respectively. Thus, $|\mathcal{N}|=n=p+q+r$.
Given an allocation $X$ for a non-degenerate instance with three types of agents, for any two agents $a_{i}$ and $a_{j}$ of the same type, we know that either $X_{a_{i}}>X_{a_{j}}$ or $X_{a_{i}}<X_{a_{j}}$. Therefore, w.l.o.g. we assume that the bundles are sorted among the agents in each group and therefore the agents $a_{1},b_{1}$ and $c_{1}$ have the least valued bundles of their respective groups. We call them the leading agents, denoted by $L=\{a_{1},b_{1},c_{1}\}$. Throughout this paper, we will maintain this property as an invariant for each EFX allocation.
@@ 2.1 Three Types of Agents
Let $X$ be an EFX allocation. For any group $T\in\{\mathbb{A},\mathbb{B},\mathbb{C}\}$ with valuation $v_{t}$, if $\exists i>j$ such that $v_{t}(X_{t}^{i})<v_{t}(X_{t}^{j})$ then we interchange the bundles of the two agents. Thus, we maintain the invariant that $v_{t}(X_{t}^{i})\geq v_{t}(X_{t}^{j})$ whenever $i>j$, for each group $T$ with valuation function $v_{t}$. In particular, the leading agent within each group has the minimum valued bundle within that group.
For an instance with three types of agents, Proposition 2 can be refined as follows.
@@
Given an instance with $3$ types of agents, and an allocation $X$, only leading agents can be the sources in the envy graph $E_{X}$. Hence $E_{X}$ has at most $3$ sources. Moreover, the following hold:
@@
    In the champion graph $M_{X}$, there must be a cycle of length at most $3$, and the cycle involves only the leading agents.
@@
To see Part 2, consider allocation of $g$ to each of the leading agents. From Proposition 4, we know that $M_{X}$ is cyclic. If a leading agent self-$g$-champions, then $M_{X}$ has a cycle of length $1$. If no leading agent self-$g$-champions, then by Part 1, no agent in any group $g$-champions the leading agent in the same group. So the leading agent of each group must be $g$-championed by a leading agent of another group, thereby creating a $2$-cycle or a $3$-cycle. ∎
@@ 2.2 Progress Measures
A common approach to proving the existence of EFX allocations is to begin with an empty, trivially EFX allocation, and iteratively extend it. To ensure termination, various progress measures have been proposed. Pareto dominance is one such well-studied measure, where each partial EFX allocation Pareto dominates the previous one. Chaudhury et al. [efx_3] introduced the lexicographic improvement to prove the existence of EFX for three agents.
To prove Theorem 1, we use a combination of Pareto dominance and a new potential function as defined below.
Definition 10 (Potential function: ).
For any, possibly partial, allocation $X$, potential of $X$ is defined to be
$\phi(X)=\min_{a_{i}\in\mathbb{A}}v_{a}(X_{a_{i}})$
@@
If $\phi(Y)>\phi(X)$, we say $Y$ $\phi$-dominates $X$. We start with a trivially EFX allocation and iteratively find new EFX allocations that either Pareto dominate or $\phi$-dominate the previous one. As Pareto dominance is a partial order on the set of allocations, and $\phi$ forms a total order, this procedure must terminate.
@@
Given a partial EFX allocation $X$, we discuss a few simple methods to obtain a $\phi$-dominating EFX allocation.
@@
Firstly, if agent $a_{1}$ is the only source in the envy graph $E_{X}$, then as a corollary of Lemma 2, we get a $\phi$-dominating EFX allocation. This is true as the bundle value of agent $a_{1}$ strictly increases.
@@
We generalize this idea to cases where there is no actual cycle among the agents, but rather among the groups. We call this the $\phi$-improving exchanges, as discussed in the following section.
2.3 $\phi$-improving exchanges
Similar to Pareto Improving Cycles in [BergerCFF22], we define $\phi$-improving exchanges for instances with three types of agents. These are not necessarily cycles among the agents, but rather cycles among the groups. Some examples of a $\phi$-improving exchanges are given in Figure 1.
Definition 11 ($\phi$-improving exchange).
[…]
(a) $\langle(a_{1},b_{i}),(b_{1},c_{j}),(c_{1},a_{1})\rangle$ is a $\phi$-improving exchange involving two compatible champion edges.
[…]
(b) $\langle(a_{1},b_{i}),(b_{1},c_{1}),(c_{1},a_{1})\rangle$ is a $\phi$-improving exchange involving one $g$-champion edge
[…]
(c) $\langle(a_{1},b_{i}),(b_{1},c_{j}),(c_{1},a_{1})\rangle$ is a $\phi$-improving exchange with only envy edges
[…]
(d) $\langle(a_{1},b_{i}),(b_{1},a_{1})\rangle$ is a $\phi$-improving exchange
Figure 1: Some examples of $\phi$-improving exchanges
Given an EFX allocation $X$, possibly partial, a $\phi$-improving exchange involving three groups is a tuple of edges either of the form $\langle(a_{1},b_{i}),(b_{1},c_{j}),(c_{1},a_{1})\rangle$ or $\langle(a_{1},c_{j}),(c_{1},b_{i}),(b_{1},a_{1})\rangle$, where $b_{i}\in\mathbb{B}$ and $c_{j}\in\mathbb{C}$, such that the following holds:
  1. (i)
  2. (ii)
    For any unallocated good $g$, the tuple can have at most one $g$-champion edge.
Similarly, a $\phi$-improving exchange involving two groups is a tuple of the form $\langle(a_{1},b_{i}),(b_{1},a_{1})\rangle$ (or $\langle(a_{1},c_{j}),(c_{1},a_{1})\rangle$) such that the above conditions hold.
To perform a $\phi$-improving exchange, within each group, the bundle that is being envied (or championed) is given to the envying (or championing) group. Within each group, the incoming bundle is then allocated to the agent who was being envied (or championed) in the exchange. If the edge is a champion edge, the $\mathsf{MES}$ bundle is transferred instead of the whole bundle.
Performing a $\phi$-exchange on an EFX allocation $X$ immediately gives a $\phi$-dominant EFX allocation $Y$. This is stated in the following Lemma 3.
Lemma 3.
A $\phi$-improving exchange on an EFX allocation $X$ gives another EFX allocation $Y$ such that $\phi(Y)>\phi(X)$.
Proof.
Consider an EFX allocation $X$. A $\phi$-improving exchange in $X$ can either involve agents from two groups, or from all the three groups.
If it involves two groups, say $\langle(a_{1},b_{i}),(b_{1},a_{1})\rangle$, we form a new allocation $Y$ by assigning $X_{a_{1}}$ to $b_{i}$ (or $\mathsf{MES}_{b}(X_{a_{1}}\cup g)$ if $(a_{1},b_{i})$ is a $g$-champion edge) and $X_{b_{i}}$ (or $\mathsf{MES}_{a}(X_{b_{i}}\cup h)$ for an $h$-champion edge) to $a_{i}$.
If it involves three groups, w.l.o.g let $\langle(a_{1},b_{i}),(b_{1},c_{j}),(c_{1},a_{1})\rangle$ be a $\phi$-improving exchange. The new allocation $Y$ is obtained by transferring the bundle of $b_{i}$ to $a_{1}$, that of $c_{j}$ to $b_{1}$, and the bundle of $a_{1}$ to $c_{1}$. If any of these are champion edges, we allocate $\mathsf{MES}$ bundles accordingly.
The bundles are then re-sorted to restore the Ordering Invariant. Let $X(A)$ be the set of bundles of agents in $\mathbb{A}$ in the allocation $X$. In allocation $Y$, the value of the minimum valued bundle in $\mathbb{A}$ increases since the new bundles in $\mathbb{A}$ i.e. $Y(A)=(X(A)\setminus\{X_{a_{1}})\}\cup\{X_{b_{i}}\}$, and $v_{a}(X_{b_{i}})>v_{a}(X_{a_{1}})$. So the new minimum valued bundle in $\mathbb{A}$ is $\min\{X_{a_{2}},X_{b_{i}}\}$.
Now we prove that $Y$ is an EFX allocation. For each agent, the bundle in $Y$ is either the same as that in $X$, or a better bundle than that in $X$. Also, the exchanged bundles are either the same bundles as in $X$ or are an $\mathsf{MES}$ of $g$ added to a bundle in $X$. Therefore, the exchanged bundles are not strongly envied by any agent. This proves that $Y$ is also an EFX allocation. ∎
2.4 Proof outline
To prove Theorem 1, we show that, for any partial EFX allocation $X$, there is another EFX allocation $Y$ such that either $Y\succ X$ or $\phi(Y)>\phi(X)$. Since the number of allocations is finite, this process must terminate with a complete EFX allocation. The proof of Theorem 1 is organized into the following case configurations, where in each configuration, we show how to construct the allocation $Y$.
  1. 1.
    Agent $a_{1}$ is a source in $E_{X}$: This is further subdivided into the following sub-cases:
    1. (a)
      Agent $a_{1}$ envies someone in $\mathbb{B}\cup\mathbb{C}$. This configuration is considered in Lemma 7 and Lemma 8 in Section 4.
    2. (b)
      Agent $a_{1}$ envies no one in $\mathbb{B}\cup\mathbb{C}$. The existence of $Y$ is proved in Lemma 9 in Section 5.
  2. 2.
    Agent $a_{1}$ is not a source in $E_{X}$: This case is considered in Section 6. Two sub-cases arise, as mentioned below:
    1. (a)
      Agent $a_{1}$ envies some agent in $\mathbb{B}\cup\mathbb{C}$. In this case, the construction of the allocation $Y$ is shown in Lemma 14.
    2. (b)
      Agent $a_{1}$ does not envy any agent in $\mathbb{B}\cup\mathbb{C}$. This case is considered in Lemma 15.
Key recurring configurations.
In proving these cases, several key configurations arise frequently, which help in constructing the desired allocation $Y$.
First, we show that when there are three sources in the envy graph $E_{X}$, we can obtain $Y$,$Y\succ X$. This is shown in Lemma 6 of Section 3.
In many scenarios, we encounter allocations such that the strong envies are confined within groups, meaning there is no inter-group strong envy. We show that such envy can be resolved without reducing the potential $\phi$. This is shown more generally in the following lemma.
Lemma 4 (Resolving internal strong envies).
Let $X$ be any allocation. Suppose an agent, say $a_{i}$, strongly envies a set $S$ of agents only within her group, i.e., $S\subset\mathbb{A}$. Then, there exists another allocation $Y$ such that the following conditions hold:
  1. (i)
    $\phi(Y)=\phi(X)$
  2. (ii)
    In $Y$, agent $a_{i}$ does not strongly envy anyone.
  3. (iii)
    In $Y$, no agent in $S$ strongly envies any agent.
Proof.
Let $a_{j}\in S$. So $a_{i}$ strongly envies $a_{j}$. Then we replace $X_{a_{j}}$ by $Y_{a_{j}}=\mathsf{MES}_{a_{i}}(X_{a_{j}})$. Clearly $X_{a_{j}}$ is not the minimum valued bundle in $\mathbb{A}$. So $\phi$ doesn’t change by this replacement. Since such a replacement is done for each $a_{j}\in S$, $a_{i}$ has no strong envy towards any agent in $S$. Since $a_{i}$ did not strongly envy anyone in other groups with respect to $X$, and since all the agents outside $S$ have the same bundles in $X$ and $Y$, $a_{i}$ does not strongly envy anyone outside $S$ either. Although agents in $S$, like $a_{j}$, got smaller bundles, $Y_{a_{j}}>_{a}X_{a_{i}}$. So $a_{j}$ does not strongly envy anyone with respect to $Y$ as $a_{i}$ does not strongly envy anyone with respect to $Y$.∎
Corollary 3.
If $X$ is an allocation where no agent strongly envies another agent from a different group, then there exists an EFX allocation $Y$,possibly partial, such that $\phi(Y)=\phi(X)$.
@@ 2.4 Proof outline
There are cases where two of the leading agents champion each other, while a third agent envies neither of them, where the following lemma applies.
@@
From Proposition 3, we know that $T_{b_{1}}^{c_{1}}>_{c_{1}}T_{c_{1}}^{b_{1}}$ and $T_{c_{1}}^{b_{1}}>_{b_{1}}T_{b_{1}}^{c_{1}}$. Therefore, both agents $b_{1}$ and $c_{1}$ are strictly better off. Now, consider the potential strong envies in $X^{\prime}$. Since the allocation $X$ was EFX, any strong envies in $X^{\prime}$ must be towards $b_{1}$ or $c_{1}$. As $T_{b_{1}}^{c_{1}}>_{c_{1}}T_{c_{1}}^{b_{1}}$, we know that
$X^{\prime}_{c_{1}}=T_{b_{1}}^{c_{1}}\cup B_{c_{1}}^{b_{1}}>T_{c_{1}}^{b_{1}}\cup B_{c_{1}}^{b_{1}}=X_{c_{1}}>T_{b_{1}}^{c_{1}}\cup B_{b_{1}}^{c_{1}}=X_{b_{1}}>T_{c_{1}}^{b_{1}}\cup B_{b_{1}}^{c_{1}}=X^{\prime}_{b_{1}}$
Therefore, no agent from group $\mathbb{C}$ envies agent $b_{1}$. Similarly, no one from $\mathbb{B}$ envies agent $c_{1}$. So, the only possible strong envies are from agents in $\mathbb{A}$ to either $2$ or $3$. We know that if any agent in $\mathbb{A}$ envies some agent $a^{*}\in\mathcal{N}$, then $a_{1}$ also envies $a^{*}$. Since agent $a_{1}$ envied neither $b_{1}$ nor $c_{1}$ in the allocation $X$, agent $a_{1}$ can envy at most one agent among $b_{1}$ and $c_{1}$ in $X^{\prime}$. Without loss of generality, assume that agent $a_{1}$ strongly envies agent $b_{1}$. Then, we replace $B_{b_{1}}^{c_{1}}$ in agent $b_{1}$’s bundle with $g$ to get the following allocation:
@@
As $T_{c_{1}}^{b_{1}}\cup g$ is a minimally envied subset, no agent strongly envies $b_{1}$. Since $T_{c_{1}}^{b_{1}}\cup g>_{b_{1}}X_{b_{1}}$, she does not envy agent $c_{1}$. Furthermore, since agent $c_{1}$ did not self $g$-champion in $X$, no agent in $\mathbb{C}$ envy agent $b_{1}$. Finally, if any agent from $\mathbb{C}\setminus{c_{1}}$ strongly envies $c_{1}$, we replace $c_{1}$’s bundle with $\mathsf{MES}_{\mathbb{C}}(Y_{c_{1}},X_{c_{2}})$, which is still better than $X_{c_{1}}$ for agent $c_{1}$. Therefore, $Y$ is an EFX allocation such that $Y(L)\succ X(L)$ and $Y(\mathcal{N}\setminus{L})=X(\mathcal{N}\setminus{L})$.∎

### v3

@@ 2 Preliminaries
Consider an allocation $X$. Suppose agent $a_{i}$ $g$-champions agent $a_{j}$ and does not $g$-champion $a_{\ell}$. Suppose agent $a_{k}$, for $k\neq i$, $g$-champions agent $a_{\ell}$. The bundle $X_{a_{j}}$ is decomposed as $T_{a_{j}}^{a_{i}}$ and $B_{a_{j}}^{a_{i}}$. Similarly, $X_{a_{\ell}}$ is decomposed into $T_{a_{\ell}}^{a_{k}}$ and $B_{a_{\ell}}^{a_{k}}$. Then $T_{a_{j}}^{a_{i}}>_{a_{i}}T_{a_{\ell}}^{a_{k}}$.
@@
Agent $a_{i}$ $g$-champions $a_{j}$ but not $a_{\ell}$. So, $T_{a_{j}}^{a_{i}}\cup g>_{a_{i}}X_{a_{i}}>_{a_{i}}T_{a_{\ell}}^{a_{k}}\cup g$, and $T_{a_{j}}^{a_{i}}>_{a_{i}}T_{a_{\ell}}^{a_{k}}$. ∎
3 Three Types of Agents
Given an instance $\langle\mathcal{N},\mathcal{M},\mathcal{V}\rangle$ of discrete fair division, we say that the agents are of three types if the set of agents can be partitioned as $\mathcal{N}=\mathbb{A}\uplus\mathbb{B}\uplus\mathbb{C}$, where agents in each of the sets $\mathbb{A}$, $\mathbb{B}$, and $\mathbb{C}$ have identical valuation functions. Let $p,q$ and $r$ be the number of agents in parts $\mathbb{A},\mathbb{B}$ and $\mathbb{C}$ respectively. Thus, $|\mathcal{N}|=n=p+q+r$. We refer to these as the three groups of agents. In particular, we denote the agents in group $\mathbb{A}$ as $\mathbb{A}=\{a_{1},a_{2},a_{3},\dots,a_{p}\}$, all of whom have the same additive valuation function $v_{a}$. Similarly, the agents in group $\mathbb{B}=\{b_{1},b_{2},\dots,b_{q}\}$ and group =$\mathbb{C}=\{c_{1},c_{2},\dots,c_{r}\}$ have identical additive valuation functions $v_{b}$ and $v_{c}$, respectively.
For concise notation, we use $S>_{a}T$ as a shorthand for $S>_{a_{i}}T$ for agents $a_{i}\in\mathbb{A}$. Similarly, we use $\geq_{a},<_{a},\leq_{a},=_{a}$ and $\mathsf{MES}_{a}(\cdot)$ to denote $\geq_{a_{i}},<_{a_{i}},\leq_{a_{i}},=_{a_{i}}$ and $\mathsf{MES}_{a_{i}}(\cdot)$ respectively. Likewise, we use subscripts $b$ and $c$ for agents in groups $\mathbb{B}$ and $\mathbb{C}$, respectively.
Consider an allocation $X$ for a non-degenerate instance with three types of agents. For any two agents $a_{i}$ and $a_{j}$ of the same type, say $\mathbb{A}$, we know that either $X_{a_{i}}>_{a}X_{a_{j}}$ or $X_{a_{i}}<_{a}X_{a_{j}}$. Therefore, w.l.o.g. we assume that the bundles are sorted among the agents in each group and therefore the agents $a_{1},b_{1}$ and $c_{1}$ have the least valued bundles in their respective groups. We call them the leading agents, denoted by $L=\{a_{1},b_{1},c_{1}\}$. Throughout this paper, we will maintain this property as an invariant for each EFX allocation.
@@ 3 Three Types of Agents
Let $X$ be an EFX allocation. For any group $\mathbb{T}\in\{\mathbb{A},\mathbb{B},\mathbb{C}\}$ with valuation $v_{t}$, if $\exists i>j$ such that $v_{t}(X_{t}^{i})<v_{t}(X_{t}^{j})$ then we interchange the bundles of the two agents. Thus, we maintain the invariant that $v_{t}(X_{t}^{i})\geq v_{t}(X_{t}^{j})$ whenever $i>j$, for each group $\mathbb{T}$ with valuation function $v_{t}$. In particular, the leading agent within each group has the minimum-valued bundle according to the valuation function of that group.
For an instance with three types of agents, we make the following observations on the structure of the envy and champion graphs.
@@
Given an instance with $3$ types of agents, and a partial allocation $X$, only leading agents can be the sources in the envy graph $E_{X}$. Hence $E_{X}$ has at most $3$ sources. Moreover, the following hold:
@@
    In the champion graph $M_{X}$ w.r.t. an unallocated good $g$, there must be a cycle of length at most $3$, and the cycle involves only the leading agents.
@@
To see Part 2, consider allocations of $g$ to each of the leading agents. From Proposition 2, we know that $M_{X}$ is cyclic. If a leading agent self-$g$-champions, then $M_{X}$ has a cycle of length $1$. If no leading agent self-$g$-champions, then by Part 1, no agent in any group $g$-champions the leading agent in the same group. So the leading agent of each group must be $g$-championed by a leading agent of another group, thereby creating a $2$-cycle or a $3$-cycle. ∎
@@ 3.1 Progress Measures
A common approach to proving the existence of EFX allocations is to begin with an empty, trivially EFX allocation, and iteratively improve it. To quantify the improvement and ensure termination, various progress measures have been proposed. Pareto dominance is one such well-studied measure, where each partial EFX allocation Pareto dominates the previous one. However, [efx_3] showed that this approach fails to guarantee EFX even for three agents. Instead, they used lexicographic improvement to prove the existence of EFX for three agents.
To prove Theorem 1, we use a combination of Pareto dominance and a new potential function defined below.
Definition 11 (Potential Function).
For any (possibly partial) allocation $X$, the potential of $X$ is defined as:
$\phi(X)=\min_{a_{i}\in\mathbb{A}}v_{a}(X_{a_{i}}).$
We mention that the choice of group $\mathbb{A}$ in the definition of $\phi$ is arbitrary. However, the structure of the proof of the main result depends on the identity of the group of agents with respect to which the potential function is defined, which, in this case is $\mathbb{A}$.
@@
If $\phi(Y)>\phi(X)$, we say $Y$ $\phi$-dominates $X$. We start with a trivial EFX allocation and iteratively find new EFX allocations that either Pareto dominate or $\phi$-dominate the preceding one. As Pareto dominance is a partial order on the set of allocations, and $\phi$ forms a total order, this procedure must terminate.
When we encounter allocations such that the strong envies are confined to agents within the same group(s), that is, there is no inter-group strong envy, such envy can be resolved without reducing the potential $\phi$. This is shown more generally in the following lemma.
Lemma 3 (Resolving internal strong envies).
Let $X$ be any allocation. Suppose an arbitrary agent, say $t\in\mathbb{T}$ for $\mathbb{T}\in\{\mathbb{A},\mathbb{B},\mathbb{C}\}$, strongly envies a set $S$ of agents only within her group, i.e., $S\subset\mathbb{T}$. Then, there exists another allocation $Y$ such that the following conditions hold:
  1. 1.
    $\phi(Y)=\phi(X)$
  2. 2.
    In $Y$, agent $t$ does not strongly envy anyone.
  3. 3.
    In $Y$, no agent in $S$ strongly envies any agent.
  4. 4.
    Bundles of agents in the other two groups remain unchanged.
Proof.
Without loss of generality, let $\mathbb{T}=\mathbb{A}$ and $t=a_{i}$. Let $a_{j}\in S$. So $a_{i}$ strongly envies $a_{j}$. Then we replace $X_{a_{j}}$ by $Y_{a_{j}}=\mathsf{MES}_{a_{i}}(X_{a_{j}})$. Clearly $X_{a_{j}}$ is not the minimum valued bundle in $\mathbb{A}$. So $\phi$ doesn’t change by this replacement. Since such a replacement is done for each $a_{j}\in S$, $a_{i}$ has no strong envy towards any agent in $S$. Since $a_{i}$ did not strongly envy anyone in other groups with respect to $X$, and since all the agents outside $S$ have the same bundles in $X$ and $Y$, $a_{i}$ does not strongly envy anyone outside $S$ either. Although agents in $S$, like $a_{j}$, got smaller bundles, $Y_{a_{j}}>_{a}X_{a_{i}}$. So $a_{j}$ does not strongly envy anyone with respect to $Y$ as $a_{i}$ does not strongly envy anyone with respect to $Y$.∎
Corollary 2.
If $X$ is an allocation in which no agent strongly envies another agent from a different group, there exists an EFX allocation $Y$, possibly partial, such that $\phi(Y)=\phi(X)$.
3.2 Pseudo-Cycles
Similar to Pareto Improving Cycles in [BergerCFF22], we define pseudo-cycles for instances with three types of agents. These are not necessarily cycles among the agents, but rather cycles among the groups. Some examples of pseudo-cycles are given in Figure 1.
Definition 12 (Pseudo-Cycle).
Given a possibly partial EFX allocation $X$, a pseudo-cycle involving three groups is a tuple of edges either of the form $\langle(a_{1},b_{i}),(b_{1},c_{j}),(c_{1},a_{1})\rangle$ or of the form $\langle(a_{1},c_{j}),(c_{1},b_{i}),(b_{1},a_{1})\rangle$, where $b_{i}\in\mathbb{B}$ and $c_{j}\in\mathbb{C}$, such that the following holds:
  1. 1.
  2. 2.
    For each unallocated good $g$, the tuple can have at most one $g$-champion edge.
Similarly, a pseudo-cycle involving two groups is a tuple of the form $\langle(a_{1},b_{i}),(b_{1},a_{1})\rangle$ (or $\langle(a_{1},c_{j}),(c_{1},a_{1})\rangle$) such that the above conditions hold.
[…]
(a) $\langle(a_{1},b_{i}),(b_{1},c_{j}),(c_{1},a_{1})\rangle$ is a pseudo-cycle with only envy edges
[…]
(b) $\langle(a_{1},b_{i}),(b_{1},a_{1})\rangle$ is a pseudo-cycle
[…]
(c) $\langle(a_{1},b_{i}),(b_{1},c_{j}),(c_{1},a_{1})\rangle$ is a pseudo cycle involving two compatible champion edges.
[…]
(d) $\langle(a_{1},b_{i}),(b_{1},c_{1}),(c_{1},a_{1})\rangle$ is a pseudo cycle involving one $g$-champion edge
Figure 1: Some examples of pseudo-cycles
3.3 Elementary Cases for Pareto and $\phi$-Improvements
@@
Given a partial EFX allocation $X$, we discuss a few simple methods to obtain a Pareto dominating or a $\phi$-dominating EFX allocation.
Elimination of a pseudo-cycle: A pseudo-cycle is eliminated by giving the envied bundle to the envying group. Thus a group that loses a bundle also receives a bundle. While the received bundle may have a lower value than the lost bundle, note that it is still higher valued than the minimum bundle in that group. Within the group, the newly received bundle is then allocated to an agent, with possible shuffling of existing bundles, so as to maintain the ordering invariant. As $a_{1}$ is always an envying and envied agent in a pseudo-cycle, the following lemma holds.
Lemma 4.
Eliminating a pseudo-cycle in an EFX allocation $X$ gives another EFX allocation $Y$ such that $\phi(Y)>\phi(X)$.
Proof.
Consider an EFX allocation $X$. A pseudo-cycle in $X$ can either involve agents from two groups, or from all the three groups.
If it involves two groups, say $\langle(a_{1},b_{i}),(b_{1},a_{1})\rangle$, we form a new allocation $Y$ by assigning $X_{a_{1}}$ to $b_{1}$ (or $\mathsf{MES}_{b}(X_{a_{1}}\cup g)$ if $(b_{1},a_{1})$ is a $g$-champion edge) and $X_{b_{i}}$ (or $\mathsf{MES}_{a}(X_{b_{i}}\cup h)$ for an $h$-champion edge) to $a_{i}$.
If it involves three groups, w.l.o.g let $\langle(a_{1},b_{i}),(b_{1},c_{j}),(c_{1},a_{1})\rangle$ be a pseudo-cycle. The new allocation $Y$ is obtained by transferring the bundle of $b_{i}$ to $a_{1}$, that of $c_{j}$ to $b_{1}$, and the bundle of $a_{1}$ to $c_{1}$. If any of these are champion edges, we allocate $\mathsf{MES}$ bundles accordingly.
The bundles are then re-sorted to restore the Ordering Invariant. Let $X(\mathbb{A})$ be the set of bundles of agents in $\mathbb{A}$ in the allocation $X$. In allocation $Y$, the value of the minimum valued bundle in $\mathbb{A}$ increases since the new bundles in $\mathbb{A}$ i.e. $Y(\mathbb{A})=(X(\mathbb{A})\setminus\{X_{a_{1}})\}\cup\{X_{b_{i}}\}$, and $v_{a}(X_{b_{i}})>v_{a}(X_{a_{1}})$. So the new minimum valued bundle in $\mathbb{A}$ is $\min\{X_{a_{2}},X_{b_{i}}\}$.
Now we prove that $Y$ is an EFX allocation. For each leading agent in $Y$, the bundle in $Y$ is either the same as that in $X$, or better than $X$. Also, the exchanged bundles are either the same bundles as in $X$ or are an $\mathsf{MES}$ bundle in $X$. Therefore, the exchanged bundles are not strongly envied by any of the leading agents. Furthermore, the agents whose utility decreased in $Y$ still have a higher utility compared to their respective leading agents. Therefore, by applying Proposition 4, none of the non-leading agents have strong envy towards any agent. This proves that $Y$ is also an EFX allocation. ∎
@@
Another way to find a $\phi$-dominating EFX allocation is when agent $a_{1}$ is the only source in the envy graph $E_{X}$. In this case, as a corollary of Lemma 2, we obtain a $\phi$-dominating EFX allocation. This holds because the bundle value of agent $a_{1}$ strictly increases.
@@
There are cases where two of the leading agents champion each other, while a third agent envies neither of them; in these cases, we can make either Pareto improvements or $\phi$-improvements, using the following lemma:
@@
From Proposition 3, we know that $T_{b_{1}}^{c_{1}}>_{c}T_{c_{1}}^{b_{1}}$ and $T_{c_{1}}^{b_{1}}>_{b}T_{b_{1}}^{c_{1}}$. Therefore, both agents $b_{1}$ and $c_{1}$ are strictly better off. Now, consider the potential strong envies in $X^{\prime}$. Since the allocation $X$ was EFX, any strong envies in $X^{\prime}$ must be towards $b_{1}$ or $c_{1}$. As $T_{b_{1}}^{c_{1}}>_{c}T_{c_{1}}^{b_{1}}$, we know that
$X^{\prime}_{c_{1}}=_{c}T_{b_{1}}^{c_{1}}\cup B_{c_{1}}^{b_{1}}>_{c}T_{c_{1}}^{b_{1}}\cup B_{c_{1}}^{b_{1}}=_{c}X_{c_{1}}>_{c}T_{b_{1}}^{c_{1}}\cup B_{b_{1}}^{c_{1}}=_{c}X_{b_{1}}>_{c}T_{c_{1}}^{b_{1}}\cup B_{b_{1}}^{c_{1}}=_{c}X^{\prime}_{b_{1}}$
Therefore, no agent from group $\mathbb{C}$ envies agent $b_{1}$. Similarly, no one from $\mathbb{B}$ envies agent $c_{1}$. So, the only possible strong envies are from agents in $\mathbb{A}$ to either $b_{1}$ or $c_{1}$. We know that if any agent in $\mathbb{A}$ envies some agent $a^{*}\in\mathcal{N}$, then $a_{1}$ also envies $a^{*}$. Since agent $a_{1}$ envied neither $b_{1}$ nor $c_{1}$ in the allocation $X$, agent $a_{1}$ can envy at most one agent among $b_{1}$ and $c_{1}$ in $X^{\prime}$. If $a_{1}$ envies neither, then we arbitrarily pick $b_{1}$. Otherwise, we pick the agent that $a_{1}$ envies, say $b_{1}$.We replace $B_{b_{1}}^{c_{1}}$ in agent $b_{1}$’s bundle with $g$ to get the following allocation:
@@
As $T_{c_{1}}^{b_{1}}\cup g$ is a minimally envied subset, no agent strongly envies $b_{1}$. Since $T_{c_{1}}^{b_{1}}\cup g>_{b}X_{b_{1}}$, she does not envy agent $c_{1}$. Furthermore, since agent $c_{1}$ did not self $g$-champion in $X$, no agent in $\mathbb{C}$ envy agent $b_{1}$. Finally, if any agent from $\mathbb{C}\setminus{c_{1}}$ strongly envies $c_{1}$, we replace $c_{1}$’s bundle with $\mathsf{MES}_{c}(Y_{c_{1}},X_{c_{2}})$, which is still better than $X_{c_{1}}$ for agent $c_{1}$. Therefore, $Y$ is an EFX allocation such that $Y(L)\succ X(L)$ and $Y(\mathcal{N}\setminus{L})=X(\mathcal{N}\setminus{L})$.∎